# Frequency-Adaptive Dilated Convolution for Semantic Segmentation

### v1

@@ Abstract
Dilated convolution, which expands the receptive field by inserting gaps between its consecutive elements, is widely employed in computer vision. In this study, we propose three strategies to improve individual phases of dilated convolution from the view of spectrum analysis. Departing from the conventional practice of fixing a global dilation rate as a hyperparameter, we introduce Frequency-Adaptive Dilated Convolution (FADC) which dynamically adjusts dilation rates spatially based on local frequency components. Subsequently, we design two plug-in modules to directly enhance effective bandwidth and receptive field size. The Adaptive Kernel (AdaKern) module decomposes convolution weights into low-frequency and high-frequency components, dynamically adjusting the ratio between these components on a per-channel basis. By increasing the high-frequency part of convolution weights, AdaKern captures more high-frequency components, thereby improving effective bandwidth. The Frequency Selection (FreqSelect) module optimally balances high- and low-frequency components in feature representations through spatially variant reweighting. It suppresses high frequencies in the background to encourage FADC to learn a larger dilation, thereby increasing the receptive field for an expanded scope. Extensive experiments on segmentation and object detection consistently validate the efficacy of our approach. The code is made publicly available at https://github.com/Linwei-Chen/FADC.
@@ 1 Introduction
Figure 2: Overview of the proposed Frequency Adaptive Dilated Convolution (FADC). It comprises three strategies, Adaptive Dilation Rate (AdaDR), Adaptive Kernel (AdaKern), and Frequency Selection (FreqSelect).
@@
Here, we introduce Frequency-Adaptive Dilated Convolution (FADC) to enhance dilated convolution through the lens of spectrum analysis. As illustrated in Figure 2, FADC comprises three key strategies, i.e., Adaptive Dilation Rate (AdaDR), Adaptive Kernel (AdaKern), and Frequency Selection (FreqSelect), aimed at enhancing the individual phases of vanilla dilated convolution. AdaDR spatially adjusts the dilation rate, AdaKern operates on the convolution kernel weights, while FreqSelect directly balances the frequency power of the input feature to encourage the expansion of the receptive field.
@@
The FreqSelect increases the receptive field size by balancing high- and low-frequency components in the feature before feeding into dilated convolution. Since convolution tends to amplify high-frequency components [40], features after dilated convolution often exhibit a higher proportion of high-frequency components. To capture these added high-frequency components, a small dilation rate $D$ will be favored for its large effective bandwidth, at the cost of compromised receptive field size. By suppressing the high-frequency power on the input features, our FreqSelect module is able to increase the respective field size. Specifically, As shown in Fig.2, FreqSelect decomposes the feature map into 4 frequency channels from low to high. Then, we spatially reweights each channel with a selection map to balance frequency power, enabling FADC to effectively learn a larger receptive field.
Our experimental results in segmentation show that our proposed method consistently brings improvements, thus validating the effectiveness of our approach. In particular, when our proposed method is applied with PIDNet, it achieves the optimal balance between inference speed and accuracy on Cityscapes, resulting in an 81.0 mIoU at 37.7 FPS. Moreover, our proposed strategy can also be integrated into deformable convolution and dilated attention, resulting in a consistent boost in performance for both segmentation and object detection tasks. Our contributions can be summarized as follows:
@@
  • We introduced Frequency-Adaptive Dilated Convolution (FADC). It adopts Adaptive Dilation Rate (AdaDR), Adaptive Kernel (AdaKern), and Frequency Selection (FreqSelect) strategies. AdaDR dynamically adjusts dilation rates in a spatially variant manner to achieve a balance between effective bandwidth and receptive field. AdaKern adaptive adjusts the kernel to fully utilize the bandwidth and FreqSelect learns a frequency-balanced feature to encourage a large receptive field.
@@ 2 Related work
Content-Adaptive Networks. The effectiveness of content-adaptive characteristics has been demonstrated by various works [14, 57, 51, 75, 8, 49]. One content-adaptive strategy involves weight adjustments, which are widely employed. Recent vision transformers [10, 35, 16] incorporate attention mechanisms to predict input-adaptive attention values. These models have achieved significant success with large receptive, but suffer from heavy computation.
@@ 3.1 Adaptive Dilation Rate
$\mathbf{\hat{D}}(p)$ can be predicted by a convolutional layer with parameters $\theta$. Particularly, we incorporate a ReLU layer to ensure non-negativity of the dilations. It aims to maximize the receptive field and minimize the lost frequency information for each pixel. For a local feature centered at $p$ with a window size of $s$, we term it as $\mathbf{X}^{(p,s)}$. Its receptive field $\text{RF}(p)=(K-1)\times\hat{\mathbf{D}}(p)+1$ is positively related to $\hat{\mathbf{D}}(p)$. The frequencies in a set $\mathcal{H}^{+}_{\mathbf{\hat{D}(p)}}$ are unable to be captured accurately. Thus, the lost frequency information can be measured by calculating the high-frequency power $\text{HP}(p)=\sum_{\mathcal{H}^{+}_{\hat{\mathbf{D}}(p)}}|\mathbf{X}^{(p,s)}_{F}(u,v)|^{2}$. Therefore, the optimization of $\theta$ can be written as
@@
Here, $\text{HP}^{+}$ and $\text{HP}^{-}$ represent pixels with the highest/lowest (e.g., 25%) high-frequency power, i.e., the brighter/darker areas in Figure 4(b), respectively. We empirically set the weight of this optimization target to 0.01 to balance it during training with task loss (e.g., pixel-wise cross-entropy for segmentation).
@@ 3.2 Adaptive Kernel
AdaDR achieves a delicate equilibrium between effective bandwidth and receptive field through the individual assignment of dilation rates to each pixel, optimizing both factors collectively. The effective bandwidth, intimately connected to the convolutional kernel’s weight, plays a pivotal role. Traditional convolutional kernels learn to capture features across diverse frequency bands, which are crucial for comprehending intricate visual patterns. However, they become static once trained. To further enhance effective bandwidth, we decompose convolutional kernel parameters into low-frequency and high-frequency components before introducing dynamic weighting to adjust the frequency response. This process only adds minor additional parameters and computational overhead. For a static convolutional kernel, its weights $\mathbf{W}$ can be decomposed as follows
@@ 4.1 Experiments Settings
Datasets and Metrics. We evaluate our methods on several challenging semantic segmentation datasets, including Cityscapes [7] and ADE20K [74]. We employ the mean Intersection over Union (mIoU) for semantic segmentation and Average Precision (AP) for object detection/instance segmentation as our evaluation metrics.
Implement details. Mask2Former [5], PIDNet [64], ResNet/HorNet+UPerNet, we keep the same setting with the original paper [5, 64, 45]. On the COCO [31] dataset, we adhere to common practices [45, 15, 60] and train object detection and instance segmentation models for 12 (1$\times$ schedule) or 36 (3$\times$ schedule) epochs. In the case of Dilated-ResNet, we substitute the dilated convolution at stage-3$\sim$4 with the proposed FADC. For PIDNet, the convolution at the bottleneck is replaced with the proposed FADC. For ResNet, we replace the convolution at stage-2$\sim$4 with the proposed FADC, while for HorNet, the convolution is replaced with the proposed FADC.
@@ 4.2 Main Results
Table 2: Quantitative comparisons of semantic segmentation models with UPerNet [63] on the ADE20K validation set.
@@
ResNet-101 [18] 85M 1025G 42.9 44.0
ResNet-101-FADC (Ours) 67M 949G 44.4 45.5
ResNet-152 [18] 85M 1025G 42.9 44.0
ResNet-152-FADC (Ours) 67M 949G 44.4 45.5
@@
Standard Semantic Segmentation. As shown in Table 1, we compared the proposed FADC with Dilated Convolution [70], Deformable Convolution (DCNv2)[76], and Adaptive Dilated Convolution (ADC)[66]. On the widely used Cityscapes dataset [7], when equipped with our FADC, the results for PSPNet, DeepLabV3+, and Mask2Former show improvements of +2.6, +1.1, and +1.2 mIoU, respectively. These enhancements outperform DCNv2 by 0.7, 0.4, and 0.2 mIoU with fewer additional computations and parameters. FADC also outperforms ADC, which adopts an adaptive dilation strategy, by 0.8 mIoU. Furthermore, as demonstrated in Table 2 using the more challenging ADE20K dataset, FADC significantly enhances the mIoU of ResNet-50 with UPerNet by 3.7, surpassing even its heavier counterpart, ResNet-101 (44.4 vs. 42.9). When applied with larger HorNet-B/L, it leads to +0.6/+0.4 gains and outperforms recent state-of-the-art methods, including Swin, ConvNeXt, RepLKNet-31L, InternImage, and DiNAT. Notably, HorNet-B-FADC exhibits superior performance and improvement (51.1 vs. 49.3, and +0.6 vs. +0.2) compared to ConvNeXt-B-dcls [23], which applies learning dilation spacing.
@@
Visualized Results. We present representative visualization results in Figure 5. The top row demonstrates that dilated convolution fails to accurately extract high-frequency information, such as the fine details of thin poles. In contrast, our proposed Frequency-Adaptive Dilated Convolution (FADC) accurately captures these details, resulting in superior predictions. In the bottom row, it is evident that dilated convolution struggles to respond uniformly to large trucks due to an insufficient receptive field to extract local information. On the other hand, FADC uniformly responds to large trucks, leading to more consistent and accurate segmentation predictions. These visualizations serve to illustrate the effectiveness of our proposed FADC in addressing the limitations of dilated convolution. Refer to supplementary for more visualization.
Table 4: Object detection results on COCO [32] dataset. All models are trained with 1$\times$ schedule. Combine the proposed AdaKern and FreqSelect strategies with deformable convolution [76].
Model: Faster-RCNN [47] Param FLOPs $\text{AP}^{b}$ $\text{AP}_{50}^{box}$ $\text{AP}_{75}^{box}$
@@
Table 5: Combine the proposed FreqSelect strategies with InternImage [60] on the ADE20K validation set.
@@
Figure 5: Visualized results on Cityscape [7]. Aliasing artifacts are evident in (a), leading to the loss of details in the representation of a thin pole and truck boundary, resulting in inferior predictions in (c). In contrast, our proposed FADC method in (b) exhibits an accurate and uniform response to both thin poles and large trucks, contributing to consistently accurate predictions in (d).
Table 6: Object detection and instance segmentation performance on the COCO [32] dataset with the Mask R-CNN [17] detector. All models are trained with 3$\times$ schedule [15, 60].
[…]
@@
Figure 6: Left: The curve illustrates the relationship between normalized high-frequency power and the predicted dilated rate. FADC, incorporating AdaDR, assigns lower dilation rates to areas with more high-frequency components, such as the boundaries of cars. Right: Mean dilated rates of layers at stages 3 to 4 in ResNet.
@@ 5 Analysis and Disccusion
Bandwidth. Measuring the bandwidth of a complex neural network is not straightforward [49], instead, we directly assess the frequency information in extracted features. As shown in Figure 3, in comparison with dilated convolution, FADC increases the power in the high frequency band of $[\frac{1}{8},\frac{1}{4})$ and $[\frac{1}{4},\frac{1}{2}]$. AdaKern further enhances the power in the frequency band $[\frac{1}{4},\frac{1}{2}]$. This indicates the extraction of more high-frequency information, demonstrating an improved bandwidth.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (61827901, 61936011, 62088101, and 62171038).

### v7

@@ Abstract
Dilated convolution, which expands the receptive field by inserting gaps between its consecutive elements, is widely employed in computer vision. In this study, we propose three strategies to improve individual phases of dilated convolution from the perspective of spectrum analysis. Departing from the conventional practice of fixing a global dilation rate as a hyperparameter, we introduce Frequency-Adaptive Dilated Convolution (FADC), which dynamically adjusts dilation rates spatially based on local frequency components. Subsequently, we design two plug-in modules to directly enhance effective bandwidth and receptive field size. The Adaptive Kernel (AdaKern) module decomposes convolution weights into low-frequency and high-frequency components, dynamically adjusting the ratio between these components on a per-channel basis. By increasing the high-frequency part of convolution weights, AdaKern captures more high-frequency components, thereby improving effective bandwidth. The Frequency Selection (FreqSelect) module optimally balances high- and low-frequency components in feature representations through spatially variant reweighting. It suppresses high frequencies in the background to encourage FADC to learn a larger dilation, thereby increasing the receptive field for an expanded scope. Extensive experiments on segmentation and object detection consistently validate the efficacy of our approach. The code is made publicly available at https://github.com/ying-fu/FADC.
@@ 1 Introduction
Figure 2: Overview of the proposed Frequency Adaptive Dilated Convolution (FADC). It comprises three strategies: Adaptive Dilation Rate (AdaDR), Adaptive Kernel (AdaKern), and Frequency Selection (FreqSelect).
@@
In this paper, we introduce Frequency-Adaptive Dilated Convolution (FADC) to enhance dilated convolution through the lens of spectrum analysis. As illustrated in Figure 2, FADC comprises three key strategies, i.e., Adaptive Dilation Rate (AdaDR), Adaptive Kernel (AdaKern), and Frequency Selection (FreqSelect), aimed at enhancing the individual phases of vanilla dilated convolution. AdaDR spatially adjusts the dilation rate, AdaKern operates on the convolution kernel weights, while FreqSelect directly balances the frequency power of the input feature to encourage the expansion of the receptive field.
@@
The FreqSelect increases the receptive field size by balancing high- and low-frequency components in the feature before feeding into dilated convolution. Since convolution tends to amplify high-frequency components [49], features after dilated convolution often exhibit a higher proportion of high-frequency components. To capture these added high-frequency components, a small dilation rate $D$ will be favored for its large effective bandwidth, at the cost of compromised receptive field size. By suppressing the high-frequency power on the input features, our FreqSelect module is able to increase the respective field size. Specifically, As shown in Figure.2, FreqSelect decomposes the feature map into 4 frequency channels from low to high. Then, we spatially reweights each channel with a selection map to balance frequency power, enabling FADC to effectively learn a larger receptive field.
Our experimental results in semantic segmentation show that our proposed method consistently brings improvements, thus validating the effectiveness of our approach. In particular, when our proposed method is applied with PIDNet, it achieves the optimal balance between inference speed and accuracy on Cityscapes, resulting in an 81.0 mIoU at 37.7 FPS. Moreover, our proposed strategy can also be integrated into deformable convolution and dilated attention, resulting in a consistent boost in performance for both segmentation and object detection tasks. Our contributions can be summarized as follows:
@@
  • We introduced Frequency-Adaptive Dilated Convolution (FADC). It adopts the Adaptive Dilation Rate (AdaDR), Adaptive Kernel (AdaKern), and Frequency Selection (FreqSelect) strategies. AdaDR dynamically adjusts dilation rates in a spatially variant manner to achieve a balance between effective bandwidth and receptive field. AdaKern adaptively adjusts the kernel to fully utilize the bandwidth, and FreqSelect learns a frequency-balanced feature to encourage a large receptive field.
@@ 2 Related work
Content-Adaptive Networks. As deep learning technology advances [74, 72, 18, 87, 85, 59, 86], the effectiveness of content-adaptive characteristics has been demonstrated by various works [21, 67, 61, 90, 13, 58]. One content-adaptive strategy involves weight adjustments, which are widely employed. Recent vision transformers [15, 43, 23] incorporate attention mechanisms to predict input-adaptive attention values. These models have achieved significant success with large receptive, but suffer from heavy computation.
@@ 3.1 Adaptive Dilation Rate
$\mathbf{\hat{D}}(p)$ can be predicted by a convolutional layer with parameters $\theta$. Particularly, we incorporate a ReLU layer to ensure the non-negativity of the dilations, and we also adopt the modulation mechanism [91]. It aims to maximize the receptive field and minimize the lost frequency information for each pixel. For a local feature centered at $p$ with a window size of $s$, we term it as $\mathbf{X}^{(p,s)}$. Its receptive field $\text{RF}(p)=(K-1)\times\hat{\mathbf{D}}(p)+1$ is positively related to $\hat{\mathbf{D}}(p)$. The frequencies in a set $\mathcal{H}^{+}_{\mathbf{\hat{D}(p)}}$ are unable to be captured accurately. Thus, the lost frequency information can be measured by calculating the high-frequency power $\text{HP}(p)=\sum_{\mathcal{H}^{+}_{\hat{\mathbf{D}}(p)}}|\mathbf{X}^{(p,s)}_{F}(u,v)|^{2}$. Therefore, the optimization of $\theta$ can be written as
@@
Here, $\text{HP}^{+}$ and $\text{HP}^{-}$ represent pixels with the highest/lowest (e.g., 25%) high-frequency power, i.e., the brighter/darker areas in Figure 4(b), respectively.
@@ 3.2 Adaptive Kernel
AdaDR achieves a delicate equilibrium between effective bandwidth and receptive field through the individual assignment of dilation rates to each pixel, optimizing both factors collectively. The effective bandwidth, intimately connected to the convolutional kernel’s weight, plays a pivotal role. Traditional convolutional kernels learn to capture features across diverse frequency bands, which are crucial for comprehending intricate visual patterns, but they become static once trained. To further enhance effective bandwidth, we decompose convolutional kernel parameters into low-frequency and high-frequency components before introducing dynamic weighting to adjust the frequency response. This process only adds minor additional parameters and computational overhead. For a static convolutional kernel, its weights $\mathbf{W}$ can be decomposed as follows
@@ 4.1 Experiments Settings
Datasets and Metrics. We evaluate our methods on several challenging semantic segmentation datasets, including Cityscapes [12] and ADE20K [89]. We employ the mean Intersection over Union (mIoU) for semantic segmentation [45, 40, 19, 4, 8] and Average Precision (AP) for object detection/instance segmentation [5, 24, 7, 6, 27] as our evaluation metrics.
Implement details. Mask2Former [10], PIDNet [76], ResNet/HorNet+UPerNet, we keep the same setting with the original paper [10, 76, 54]. On the COCO [38] dataset, we adhere to common practices [54, 22, 70] and train object detection and instance segmentation models for 12 (1$\times$ schedule) or 36 (3$\times$ schedule) epochs. In the case of Dilated-ResNet, we substitute the dilated convolution at stage-3$\sim$4 with the proposed FADC. For PIDNet, the convolution at the bottleneck is replaced with the proposed FADC. For ResNet, we replace the convolution at stage-2$\sim$4 with the proposed FADC.
@@ 4.2 Main Results
Table 2: Quantitative comparisons on semantic segmentation tasks with UPerNet [75] on the ADE20K validation set.
@@
Standard Semantic Segmentation. As shown in Table 1, we compared the proposed FADC with Dilated Convolution [82], Deformable Convolution (DCNv2)[91], and Adaptive Dilated Convolution (ADC)[78]. On the widely used Cityscapes dataset [12], when equipped with our FADC, the results for PSPNet, DeepLabV3+, and Mask2Former show improvements of +2.6, +1.1, and +1.2 mIoU, respectively. These enhancements outperform DCNv2 by 0.7, 0.4, and 0.2 mIoU with fewer additional computations and parameters. FADC also outperforms ADC, which adopts an adaptive dilation strategy, by 0.8 mIoU. Furthermore, as demonstrated in Table 2 using the more challenging ADE20K dataset, FADC significantly enhances the mIoU of ResNet-50 with UPerNet by 3.7, surpassing even its heavier counterpart, ResNet-101 (44.4 vs. 42.9). When applied with larger HorNet-B, it leads to +0.6 gains and outperforms recent state-of-the-art methods, including Swin, ConvNeXt, RepLKNet-31L, InternImage, and DiNAT. Notably, HorNet-B-FADC exhibits superior performance and improvement (51.1 vs. 49.3, and +0.6 vs. +0.2) compared to ConvNeXt-B-dcls [31], which applies learning dilation spacing.
@@
Visualized Results. We present representative visualization results in Figure 5. The top row demonstrates that dilated convolution fails to accurately extract high-frequency information, such as the fine details of thin poles. In contrast, our proposed Frequency-Adaptive Dilated Convolution (FADC) accurately captures these details, resulting in superior predictions. In the bottom row, it is evident that dilated convolution struggles to respond uniformly to large trucks due to an insufficient receptive field to extract local information. On the other hand, FADC uniformly responds to large trucks, leading to more consistent and accurate segmentation predictions. These visualizations serve to illustrate the effectiveness of our proposed FADC in addressing the limitations of dilated convolution.
Table 4: Combining the proposed AdaKern and FreqSelect strategies with DCNv2 [91] on the object detection task. All models are trained with a 1$\times$ schedule on the COCO dataset [39].
Model: Faster-RCNN [56] Param FLOPs $\text{AP}^{box}$ $\text{AP}_{50}^{box}$ $\text{AP}_{75}^{box}$
@@
Table 5: Combining the proposed FreqSelect strategies with InternImage [70] on the ADE20K validation set.
@@
Figure 5: Visualized results on Cityscape [12]. Aliasing artifacts are evident in (a), resulting in the loss of details in the representation of thin poles and truck boundaries, which leads to inferior predictions in (c). In contrast, our proposed FADC method in (b) demonstrates an accurate and uniform response to both thin poles and large trucks, thereby contributing to consistently accurate predictions in (d).
Table 6: Object detection and instance segmentation performance on the COCO dataset [39] with the Mask R-CNN detector [24]. All models are trained with a 3$\times$ schedule [22, 70].
Model Params FLOPs APbox APmask
@@
Figure 6: Left: The curve illustrates the relationship between normalized high-frequency power and the predicted dilation rate. FADC, incorporating AdaDR, assigns lower dilation rates to areas with more high-frequency components, such as the boundaries of cars. Right: Mean dilation rates at stages 3 and 4 in ResNet.
@@ 5 Analysis and Disccusion
Bandwidth. Measuring the bandwidth of a complex model is not straightforward [58], instead, we directly assess the frequency information in extracted features. In Figure 3, in comparison with dilated convolution, FADC increases the power in the high-frequency band of $[\frac{1}{8},\frac{1}{4})$ and $[\frac{1}{4},\frac{1}{2}]$. AdaKern further enhances the power in the frequency band $[\frac{1}{4},\frac{1}{2}]$. This indicates the extraction of more high-frequency information, demonstrating an improved bandwidth.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (62331006, 62171038, and 62088101), the R&D Program of Beijing Municipal Education Commission (KZ202211417048), the Fundamental Research Funds for the Central Universities, and JST Moonshot R&D Grant Number JPMJMS2011, Japan.